OpenStudy (anonymous):

If x be nearly equal to 1, show that

OpenStudy (anonymous):

\[\frac{ mx ^{n}-nx ^{m} }{ x ^{n}-x ^{m} }=\frac{ 1 }{ 1-x }\]

OpenStudy (anonymous):

Prove using only a single side

OpenStudy (anonymous):

\[x \approx 1\] Let \[x=1+h\]

OpenStudy (anonymous):

Where h be a small a small change , so small that h & higher power can be neglected

OpenStudy (anonymous):

I am stuck in the ending so plz help

Parth (parthkohli):

The denominator tends to zero.

OpenStudy (anonymous):

i didn't understand

OpenStudy (anonymous):

wait i am posting a type of this question that i did myself

Parth (parthkohli):

How do you define \(x \approx 1\)?

OpenStudy (anonymous):

wait @ParthKohli

OpenStudy (anonymous):

x is approximately equal to 1

Parth (parthkohli):

How do you define "approximate"? Do you have an epsilon?

OpenStudy (anonymous):

the question says x is nearly equal to 1 which means approx. equal to 1 so we consider h as a very small value

Parth (parthkohli):

Define "small"

Parth (parthkohli):

are you from India ?

OpenStudy (anonymous):

Do you know the geometric series?

OpenStudy (anonymous):

i am from Pakistan studying in australia

OpenStudy (anonymous):

\[\large \sum_{k=1}^{n-1}x^k =\frac{1-x^n}{1-x} \]

OpenStudy (anonymous):

Might help, not sure.

OpenStudy (anonymous):

can u explain the way i have done in prev. question that I've posted

OpenStudy (anonymous):

look i am showing i got stuck

Parth (parthkohli):

Also,\[\dfrac{a}{1 - r}\]is the sum of an infinite geometric sequence when \(|r|<1\)

Parth (parthkohli):

I don't know if that helps either.

OpenStudy (anonymous):

Using L.H.S \[\frac{ m(1+h)^n - n(1+h)^m }{ (1+h)^n - (1+h)^m }\]

OpenStudy (anonymous):

\[\frac{ m(1+nh+...)-n(1+mh+...) }{ (1+nh+...)-(1+mh+...) }\] Neglecting terms containing high powers \[\frac{ m(1+nh)-n(1+mh) }{ (1+nh)-(1+mh) }\]

OpenStudy (anonymous):

\[\frac{ m+mnh-n-mnh }{ 1+nh-1-mh }\]

OpenStudy (anonymous):

now from here what to do to prove \[\frac{ 1 }{ 1-x }\]

OpenStudy (anonymous):

remember x=1+h

OpenStudy (anonymous):

i got stuck at a point where i got -1/h

OpenStudy (anonymous):

plz help

OpenStudy (shubhamsrg):

h=x-1 -_-

OpenStudy (anonymous):

oh my... 0_0 why didn't this come to my mind

OpenStudy (anonymous):

thank u very much @shubhamsrg

OpenStudy (shubhamsrg):

:P

OpenStudy (shubhamsrg):

This is quite better than the proof I was typing. hmm..

OpenStudy (anonymous):

\[\frac{ m-n }{ nh-mh }\]

OpenStudy (anonymous):

-1/h

OpenStudy (anonymous):

\[-\frac{ 1 }{ 1-x }\]

OpenStudy (anonymous):

negative sign?? :/

OpenStudy (anonymous):

@shubhamsrg

OpenStudy (shubhamsrg):

h=x-1 and not 1-x

Parth (parthkohli):

\[-\dfrac{1}{x - 1}\]

Parth (parthkohli):

Oh, already posted. I shouldn't keep looking at my keyboard all the time >_<

OpenStudy (anonymous):

ok ok

OpenStudy (anonymous):

thanks to all :)